OpenStudy (anonymous):

Find 17th term of the sequence 5,5,11,29,65.....is?

OpenStudy (ybarrap):

Let's find the pattern to differences 5-5=0 11-5=6 29-11=18 65-29=36 and so on sequence is 0,6,18,36,... so pattern is 1*3*0,2*3*1,3*3*2,4*3*3...(n+1)*3*n,... (n+1)3n=3n^2 + 3n a0 = 5 a1 = a0 + 3*1^2 + 3*1=5 + 3+3 = 11 a2 = a1 + 3*2^2 + 3*2 = 11+3*4+6=29 a3 = a2 + 3*3^2 + 3*3 = 29 + 27+9=65 ... an = a(n-1) + 3*n^2 + 3n ... 17th term is a16 = a15 + 3*16^2 + 3*16 = a15 + 816 The terms a4 through a15 can be determined easily using a spreadsheet program like excel or by hand.